The set S consists of all integers from 1 to 2007 inclusive. For how many

elements n in S is f(n)=2n³+n²-n-2 over n²-2 an integer?


1
Expert's answer
2022-06-08T13:42:54-0400
"\dfrac{2x^3+x^2-x-2}{x^2-2}"

"=\dfrac{2x(x^2-2)+(x^2-2)+3x}{x^2-2}"

"=2x+1+\dfrac{3x}{x^2-2}"

Given "n\in \Z^+"

"n=1, 2(1)+1+\dfrac{3(1)}{(1)^2-2}=0, 0\in \Z"



"n\ge 2, \dfrac{3x}{x^2-2}\ge1"

"\dfrac{x^2-3x-2}{x^2-2}\le0"

"\dfrac{x^2-3x-2}{x^2-2}\le0"

"x^2-3x-2=0"

"x_1=\dfrac{3-\sqrt{17}}{2}, x_2=\dfrac{3+\sqrt{17}}{2}"


"x\in(-\sqrt{2}, \dfrac{3-\sqrt{17}}{2}]\cup (\sqrt{2}, \dfrac{3+\sqrt{17}}{2}]"

Since "n" is positive integer, we consider "n=2,3."


"n=2,\dfrac{3n}{n^2-2}=\dfrac{3(2)}{(2)^2-2}=3, integer"

"n=3,\dfrac{3n}{n^2-2}=\dfrac{3(3)}{(3)^2-2}=\dfrac{9}{7}, is\ not \ integer"

"n=1,n=3."